1. In a random sample of 200 persons of a town, 120 are found to be tea drinkers. In a random sample of 500 persons of another town, 240 are found to be tea drinkers. Is the proportion of tea drinkers in the two towns equal? Use 0.01 level of significance.
Chance of boy or girl by paternal parentage?
Comprehensive information is not available. The father's parentage is not recorded for all births - there are enough missing values to cast doubt on the accuracy of the results.
Why is a z-score range -3 to plus 3?
Strictly speaking, it is not, but the range -3 to +3 accounts for 99.73% of the total probability.
Suppose you were to roll a dicehow many different outcomes could occur as a result?
If you roll one cube, there are 6 possible outcomes.
If you roll two cubes, there are 36 possible outcomes. But if you can't tell
one cube from the other, then 1 - 2 looks the same to you as 2 - 1, so
there are only 18 different unique outcomes.
What is the conditional expectation function?
It is the integral (or sum) of the joint probability distribution function of the two events, integrated over the domain in which the condition is met.
You might need to indicate what the disorder is.
The number of permutations of the letter SUM is 3 factorial, or 6. Since none of those letters are repeated, the issue of repetition is not a factor. Perhaps the questioner meant to use a different word. If so, please restate the question.
a process of systematically selecting representation elements of a population
How many different arrangements are possible from the letters of the word calculator?
10!/(2!2!2!)=Answer
10 being 10 letters
There are "2" c's, a's, and l's, so you must divide those factorials with the total number.
Answer = 7257600
Is it possible to toss a coin 20 times and have it land heads-up 20 times is this likely to happen?
It is possible. It is certain with a double headed coin. With a fair coin the event has a probability of 1 in 1048576 or approximately 1 in a million.
What is a model in which each outcome has an equal probability of occurring?
A uniform distribution.
A uniform distribution.
A uniform distribution.
A uniform distribution.
What are the Guidelines for selecting clothing?
When selecting clothing, consider the occasion, comfort, and fit. Choose fabrics that suit the weather and your lifestyle, ensuring they are breathable and appropriate for the activity. Pay attention to color and style to reflect your personal taste and body shape. Lastly, prioritize quality over quantity for durability and sustainability.
What is probability of picking m in the word mathematics?
There are 11 letters and the m appears twice, so it is 2 in 11.
What are the differences between discrete and continuous distribution?
discrete distribution is the distribution that can use the value of a whole number only while continuous distribution is the distribution that can assume any value between two numbers.
What is the probability that the maximum of the two rolls is greater than or equal to 5?
I Assume you mean two separate dice rolls.
There are 2 possible chances to get 5 or higher in the first roll: 5 or 6. So, if you roll one of those, the second roll is irrelevant. So, thus far, we have a (2/6)*(6/6) chance = 12/36. If you miss it the first roll, meaning you rolled a 1-4, then the second roll has to be a 5 or 6. So there is a 4/6 chance of rolling 1-4 and a 2/6 chance of then rolling a 5 or 6. (4/6)*(2/6) = 8/36
So your total probability is 12/36 + 8/36 = 20/36 = 5/9
I bet you thought it was going to be 2/3 :)